In addition to charging an hourly rate, a repairman charges a fixed fee for making a service call. If a 2 hour service call costs 88$ and a 4-hour service call costs $136, what are the repairman's (a)fixed fee and (b) hourly rate?
Answer by checkley71(8403)

X+2Y=88
X+4Y=136 NOW SUBTRACT THESE EQUATIONS
---------------------------------
2Y-4Y=88-136
-2Y=-48
Y=-48/-2
Y=$24 IS THE HOURLY RATE.
X+2*24=88
x+48=88
x=88-48
x=$40 is the fixed fee.
proof
40+4*24=136
40+96=136
136=136